OpenStudy (anonymous):

2x-y=8 5x-2y=21 What is the ordered pair

OpenStudy (texaschic101):

do you know the substitution method ?

OpenStudy (texaschic101):

2x - y = 8 --> -y = 8 - 2x --> y = 2x - 8 Now sub 2x - 8 in for y in the other equation... 5x - 2y = 21 5x - 2(2x - 8) = 21 (distribute through the parenthesis) solve for x. Once x is found, sub that back into either of the original equations and solve for y. You can do this :)